Please help me with this problem using the pythagrean theorum. One leg of a right triangle is 8 inches long. The hypotenuse is 2 inches longer than the other leg. What is the length of the hypotenuse.
Answer by suresh(20)

By pythagrean theorem,
(hupotenuse)^2 = sum of squares of the legs
Given
leg1=8
We have to find leg2.
Let x be the length of leg2.
Given that hypotenuse(h) is 2 inches longer then x.
So h=x+2
therefore,h^2=8^2+x^2
(x+2)^2=64+x^2
x^2+4x+4=64+x^2
4x+4=64 (cancellation law)
4x=60
x=15
Hence the length of other side is 15 inches